Calculate the inverse Laplace transform of the function :

F(s) = (s+1)/ (s³+s²-6s)


1
Expert's answer
2021-01-05T13:11:50-0500

"F(s)=\frac{s+1}{s^3+s^2-6s}\\
\therefore F(s)\\
=\frac{s+1}{s^3+s^2-6s}\\
=\frac{(s+1)}{s(s-2)(s+3)}\\
=\frac{3}{10}\times \frac{1}{s-2}-\frac{1}{6}\times \frac{1}{s}-\frac{2}{15}\times \frac{1}{s+3}"

"So,"

"\mathcal{L}^{-1}\{F(s)\}
\\=\mathcal{L}^{-1}\{\frac{3}{10}\times \frac{1}{s-2}-\frac{1}{6}\times \frac{1}{s}-\frac{2}{15}\times \frac{1}{s+3}\}
\\=\frac{3}{10}\mathcal{L}^{-1}\left( \frac{1}{s-2}\right)-\frac{1}{6}\mathcal{L}^{-1}\left( \frac{1}{s}\right)-\frac{2}{15}\mathcal{L}^{-1}\left( \frac{1}{s+3}\right)\\
=\frac{3}{10}e^{2t}-\frac{1}{6}\times1-\frac{2}{15}e^{-3t}\\
=\frac{3}{10}e^{2t}-\frac{2}{15}e^{-3t}-\frac{1}{6}"


"\pmb\therefore" Inverse laplace of of "\pmb{F(s) :=\frac{3}{10}e^{2t}-\frac{2}{15}e^{-3t}-\frac{1}{6}}"